A survey worker uses a 4.5 ft stand to site a spot on a road that is 65 feet away. What is the grade of the road? Round your answer to the nearest whole percent.


1
Expert's answer
2021-10-26T16:25:32-0400

Given:-

Stand Height (h) = 4.5 ft

Road Distance (d)= 65 ft


To find:-

Grade of road(Q)= ?

"sin(Q)=\frac{h}{d}"

"Q= sin^{-1}( \frac{h}{d})"

"Q= sin^{-1}( \frac{4.5}{65})"

"Q= 3.69 \,degrees"